amanda can read her book at a rate of 7/8 of a chapter per 2/5 of an hour ? how many chapters can she read in one hour


1
Expert's answer
2021-02-24T05:58:32-0500
"\dfrac{7/8}{2/5}=\dfrac{7}{8}\cdot\dfrac{5}{2}=\dfrac{35}{16}"

Amanda can read 2 and 3/16 of a chapter in one hour.